OpenStudy (anonymous):

hello can someone explain to me how to do this ???????????????????????????????????? thank you :) i will give you a medal and become your fan : ) <3 BC parallel to DE. BC = 6, DE = 9 Which of the following could be the lengths of AB and BD respectively? ABC = ADE A.2 and 3 B.3 and 4.5 C.4 and 2 D.4 and 8

OpenStudy (anonymous):

https://media.glynlyon.com/g_geo_2013/5/groupi19.gif

hero (hero):

So basically, the given information tells us the we are dealing with similar triangles. Which means, that corresponding sides have the same ratio. In other words, the ratio of BC to DE is the same as the ratio of AB to AD \[\frac{BC}{DE} = \frac{AB}{AD}\]

hero (hero):

You are already given the values of BC and DE: \[\frac{6}{9} = \frac{AB}{AD}\] All you have to do is just insert the possible values for AB and AD both sides equal.

hero (hero):

All you have to do is just insert the possible values for AB and AD until* both sides equal.

OpenStudy (anonymous):

OH ! I UNDERSTAND ! thank you for explaining so clearly !

OpenStudy (anonymous):

so its b yay thanks

hero (hero):

Why do you say it is B?

OpenStudy (anonymous):

because 6/9=3/4.5?

hero (hero):

Why did you skip A?

OpenStudy (anonymous):

i didint is a the correct answer?

OpenStudy (anonymous):

@Hero

hero (hero):

Put it this way, you mentioned that B might be the answer so, explain how you arrived at B.